find the period for f(x)= -9 sin 5x
Answer by lwsshak3(11628)

find the period for f(x)= -9 sin 5x
Equation for sin function: y=Asin(Bx-C), A=amplitude, period=2π/B, phase shift=C/B
For given sin function:
B=5
period=2π/B=2π/5